# Xuezhang Chen  and Shihong Zhang

### v1

@@ 1 Introduction
We now pause to comment on the decay assumptions at infinity of initial data that $u(x,0)=O(|x|^{-c})$ as $|x|\to\infty$ in Theorems 1.6-1.8. For critical exponents $p_{i}^{\ast}$, such assumptions are to ensure that involved Poisson integrals are well-defined. As a result, this deduces that $c>\frac{3(n-3)}{n+3}$ in Theorems 1.6 and 1.8; $c>\frac{2(n-3)}{n+1}$ in Theorem 1.7. As will show in the proof of Theorem 1.9, we know $c=n-3-\alpha$, then the assumption (1.10) leads to $\alpha<\frac{n(n-3)}{n+1}$ and $\alpha<\frac{(n-1)(n-3)}{n+1}$, respectively. Another interpretation of assumption of $\alpha$ is as follows: If $\mathscr{B}^{3}_{3}(U)=\frac{n-3}{2}\tilde{T}^{3}_{3}U^{\frac{n+3}{n-3}}\in L^{1}(\mathbb{S}^{n})$ for a smooth conformal metric $\tilde{g}=U^{4/(n-3)}|\mathrm{d}\xi|^{2}$, then near the singular point $-\mathbf{e}_{n+1}$,
which requires $0<\alpha<\frac{n(n-3)}{n+3}$. Intuitively, $\mathscr{B}^{3}_{3}(U)\in L^{1}(\mathbb{S}^{n})$ means $U\in W^{3,1}(\mathbb{S}^{n})\hookrightarrow W^{2,\frac{n}{n-1}}(\mathbb{S}^{n})$. This roughly gives $\mathscr{B}^{3}_{2}(U)\in L^{\frac{n}{n-1}}(\mathbb{S}^{n})$, and also
which requires $\alpha<\frac{(n-1)(n-3)}{n+1}$.
In Section 6, we give the proof of Theorem 1.9, and as applications we also establish two sharp geometric inequalities over some $Q$-curvature flat conformal class of $(\mathbb{B}^{n+1},|\mathrm{d}\xi|^{2})$ for $n\geq 4$, including a sharp inequality of isoperimetric type.
@@ 2 Preliminaries
We give asymptotic behavior of two prototypes of singular integrals, coming from biharmonic Poisson integrals on $\mathbb{R}_{+}^{n+1}$.
@@
Suppose $f\in L_{\mathrm{loc}}^{1}(\mathbb{R}^{n})$ with $f(x)=O(|x|^{-a})$ as $|x|\to\infty$ for some $a\in\mathbb{R}_{+}$. Assume $20<\alpha<n/$ 2 and $a+2\alpha>n$. We define
@@
$\left|v(X)\right|\lesssim$$\begin{cases}|X|^{n-2\alpha-\min\{a,n\}}\quad&\mathrm{if~{}~{}}a\neq n\\ |X|^{-2\alpha}\log|X|&\mathrm{if~{}~{}}a=n\end{cases},\qquad\quad\mathrm{~{}~{}as~{}~{}}\quad|X|\to\infty.$
@@
$A_{1}=$$\left\{y||y|<|x|/2\right\},\qquad\qquad\qquad\qquad A_{2}=\left\{y||x-y|<|x|/2\right\},$
$A_{3}=$$\left\{y||y|\geq|x|/2,|x-y|\geq|x|/2\right\}.$
@@
$|v(x,t)|\lesssim t^{-2\alpha}\int_{A_{1}}|f(y)|\mathrm{d}y\lesssim$$\begin{cases}t^{-2\alpha}|x|^{n-\min\{a,n\}}&\mathrm{if~{}~{}}a\neq n\\ t^{-2\alpha}\log(|x|+1)&\mathrm{if~{}~{}}a=n\end{cases}$
@@
In $A_{1}$, we have $\frac{|x|}{2}\leq|x-y|\leq\frac{3|x|}{2}$ and $t^{2}+|x-y|^{2}\sim|x|^{2}$, then,
@@
$A_{21}=A_{2}\cap\{y||x-y|>t\}\qquad\mathrm{and}\qquad A_{22}=A_{2}\cap\{y||x-y|\leq t\}.$
@@
Suppose $f\in L_{\mathrm{loc}}^{1}(\mathbb{R}^{n})$ with the decay that $f(x)=O(|x|^{-b})$ as $|x|\to\infty$ for some positive constant $b$. We define
$w(x,t)=\int_{\mathbb{R}^{n}}\frac{t^{\beta}}{(t^{2}+|x-y|^{2})^{\frac{n+\beta}{2}}}f(y)\mathrm{d}y,\quad\beta\in\mathbb{R}_{+}.$
@@
$\left|w(x,t)\right|\lesssim$$\begin{cases}|X|^{-\min\{b,n\}}\qquad&\mathrm{if~{}~{}}b\neq n\\ |X|^{-n}\log|X|&\mathrm{if~{}~{}}b=n\end{cases},\qquad\quad\mathrm{~{}~{}as~{}~{}}\quad|X|\to\infty.$
@@
For $r\in\mathbb{R}_{+}$, let $B_{r}^{+}=B_{r}(0)\cap\mathbb{R}^{n+1}_{+}$ and $\partial^{+}B_{r}^{+}=\partial B_{r}^{+}\cap\mathbb{R}^{n+1}_{+}$. Let $P_{k}(x)$ denote a polynomial with degree $k$ in the variable $x$, where $k\in\mathbb{N}$. Given $u\in C^{\infty}(\overline{\mathbb{R}^{n+1}_{+}})$, we introduce a Kelvin transform for $\Delta^{2}$ by
$u^{\ast}(X)=|X|^{3-n}u\left(\frac{X}{|X|^{2}}\right),\qquad X\in\overline{\mathbb{R}_{+}^{n+1}}\backslash\{0\}.$
For our purpose, we need to refine Sun-Xiong’s results [30, Proposition 3.1 and Theorem 3.2] on homogeneous biharmonic boundary value problems.
@@
$\begin{cases}\displaystyle\Delta^{2}u=0&\mathrm{~{}~{}in~{}~{}}\quad\mathbb{R}_{+}^{n+1},\\ \displaystyle\partial_{t}u(x,0)=0&\mathrm{~{}~{}on~{}~{}}\quad\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\partial_{t}\Delta u(x,0)=0\qquad\qquad&\mathrm{~{}~{}on~{}~{}}\quad\partial\mathbb{R}^{n+1}_{+}.\end{cases}$ (2.1)
@@
It is notable that the nonnegativity of solutions is removed compared with that of Sun-Xiong [30].
@@
First, rewrite the boundary conditions as $\mathscr{B}_{1}^{3}u=0,\mathscr{B}_{3}^{3}u=0$, then it follows from conformal covariance that $u^{\ast}$ satisfies the biharmonic boundary value problem (2.1) in $\overline{\mathbb{R}_{+}^{n+1}}\backslash\{0\}$.
@@
Notice that $\eta_{\varepsilon}\in C^{3,1}(\mathbb{R}^{n+1}_{+})$ and $u^{*}$ is biharmonic on +$\overline{\mathbb{R}^{n+1}_{+}}\backslash\{0\}$, then integrating by parts gives
@@
For each $r\in\mathbb{R}_{+}$, we set $v_{r}(X)=u^{*}(rX)$, which is biharmonic in +$\overline{\mathbb{R}^{n+1}_{+}}\backslash\{0\}$. Now we extend $v_{r}$ by an even reflection to $\mathbb{R}^{n+1}\backslash\{0\}$ and apply the local estimates for biharmonic functions in [27, Proposition 4] to obtain
@@
so $\tilde{u}\in C^{4}(\mathbb{R}^{n+1})$ and is biharmonic in $\mathbb{R}^{n+1}$. By [27, Proposition 4] we obtain
@@
If we relax the above assumption (2.2) to $u(X)\geq-C|X|^{\delta}$ for some constant $40<\delta<$ 4, which ensures that $u^{\ast}_{-}\in L^{1}(B_{1}^{+})$, then we apply the argument above to show that
@@
$\begin{cases}\displaystyle\Delta^{2}u=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1},\\ \displaystyle u(x,0)=0\qquad\qquad\qquad\qquad\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\Delta u(x,0)=0\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+}.\end{cases}$ (2.4)
@@
$u(X)\geq-C|X|^{-\delta}\qquad\mathrm{as}\quad|X|\to\infty.$ (2.5)
@@
Similarly, the boundary conditions become $\mathscr{B}_{0}^{3}u=0,\mathscr{B}_{3}^{3}u=0$ on $\partial\mathbb{R}_{+}^{n+1}$. By conformal covariance, $u^{\ast}$ satisfies the biharmonic boundary value problem (2.4) on +$\overline{\mathbb{R}^{n+1}_{+}}\backslash\{0\}$. Following the same lines as before, we have $u^{\ast}\in L^{1}(B_{1}^{+})$ and $u(x,t)=o(|X|^{4})$ as $|X|\to\infty$. Finally, we extend $u$ by an odd reflection
$\tilde{u}(x,t)=\begin{cases}\displaystyle u(x,t)\qquad&\mathrm{for}\qquad t\geq 0,\\ \displaystyle-u(x,-t)\qquad&\mathrm{for}\qquad t<0.\end{cases}$
A similar argument as in Lemma 2.3 combined with boundary conditions shows
@@
If we relax the above assumption (2.5) to $u(X)\geq-C|X|^{\delta}$ as $|X|\to\infty$ for some $30<\delta<$ 3, then the same result in Lemma 2.4 still holds.
@@ 3 Biharmonic Poisson kernel
Our purpose is to derive explicit formulae of biharmonic Poisson kernels for $\mathbb{R}_{+}^{n+1}$ and $\mathbb{B}^{n+1}$.
3.1 Noncritical dimension: The upper half-space
We start with the standard Poisson kernel for $-\Delta$ in $\mathbb{R}_{+}^{n+1}$:
@@ 3.1 Noncritical dimension: The upper half-space
Suppose $f\in C^{2}(\mathbb{R}^{n})\cap L^{\infty}(\mathbb{R}^{n})$ and $\alpha>1$, then for every fixed $x\in\mathbb{R}^{n}$,
@@
It is not hard to see that the above result is true provided both $f$ and $|\nabla f|$ are bounded.
@@
Therefore, with the above choice of $\sigma$ we combine these estimates to conclude that for $t\ll 1$,
@@
solves $\Delta u=0$ in $\mathbb{R}_{+}^{n+1}$ and $u=f(x)$ on $\partial\mathbb{R}_{+}^{n+1}$. However, for each $x\notin\mathrm{supp}(f)$, the Hopf boundary point lemma yields $\partial_{t}u(x,0)>0$.
Proof of Theorem 1.1. It suffices to show that fix $i,k\in\{0,1,2,3\}$, we define
@@
Case 1. Type $(\mathscr{B}^{3}_{0},\mathscr{B}^{3}_{1})$ boundary condition.
@@
Case 2. Type $(\mathscr{B}^{3}_{0},\mathscr{B}^{3}_{2})$ boundary condition.
@@
$\partial_{t}\Delta|X|^{3-n}=$ $(3-n)\Delta(t|X|^{1-n})=2(n-1)(n-3)t|X|^{-1-n};$
@@
Case 4. Type $(\mathscr{B}^{3}_{2},\mathscr{B}^{3}_{3})$ boundary condition.
@@
3.2 Noncritical dimension: The unit ball
Our purpose is to give explicit representation formulae of biharmonic Poisson kernels for $\mathbb{B}^{n+1}$ with the help of conformal equivalence between $\mathbb{B}^{n+1}$ and $\mathbb{R}_{+}^{n+1}$,

### v2

@@ 1 Introduction
A geometric interpretation of Corollary 1.1 is that, Let $g$ be a $Q$-flat conformal metric on $(\mathbb{B}^{n+1},|\mathrm{d}\xi|^{2})$ for $n\geq 4$. Assume that $\big{(}(T^{3}_{i})_{g},(T^{3}_{j})_{g}\big{)}=\big{(}(T^{3}_{i})_{\mathbb{S}^{n}},(T^{3}_{j})_{\mathbb{S}^{n}}\big{)}$ for $i,j\in\{1,2,3\},i<j$. Then there exists a conformal diffeomorphism $\Psi$ on $\overline{\mathbb{B}^{n+1}}$ such that $g=\Psi^{\ast}(|\mathrm{d}\xi|^{2})$.
In Section 6, we give the proof of Theorem 1.7. As applications we establish two sharp geometric inequalities over certain $Q$-flat conformal class of $(\mathbb{B}^{n+1},|\mathrm{d}\xi|^{2})$ for $n\geq 4$, including a sharp fourth-order isoperimetric inequality.
In Section 7, we raise a rigidity conjecture for a geodesic ball in $\mathbb{S}^{n+1}$, which is a natural extension of Corollary 1.1, and could be reduced to the classification of positive solutions to the constant $Q$-curvature and constant $T$-curvature equation in $\mathbb{B}^{n+1}$. We prove a uniqueness result of positive solutions to a fourth-order geometric ODE in Appendix B. This allows us to answer this conjecture affirmatively for radial solutions. In Appendix A, we give an alternative proof of Proposition 3.1.
@@ 2 Preliminaries
This section is devote to building some technical lemmas and giving asymptotic behavior of two prototypes of singular integrals, which are two variants of biharmonic Poisson integrals for $\mathbb{R}_{+}^{n+1}$.
@@
Suppose $f\in L_{\mathrm{loc}}^{1}(\mathbb{R}^{n})$ with $f(x)=O(|x|^{-a})$ as $|x|\to\infty$ for some $a\in\mathbb{R}_{+}$. Assume that $20<\alpha<n/$ 2 and $a+2\alpha>n$. We define
@@
$\left|v(X)\right|\lesssim$$\begin{cases}|X|^{n-2\alpha-\min\{a,n\}}\quad&\mathrm{if~{}~{}}a\neq n\\ |X|^{-2\alpha}\log|X|&\mathrm{if~{}~{}}a=n\end{cases}\qquad\quad\mathrm{~{}~{}as~{}~{}}\quad|X|\to\infty.$
@@
$A_{1}=$$\left\{y;|y|<|x|/2\right\},\qquad\qquad\qquad\qquad A_{2}=\left\{y;|x-y|<|x|/2\right\},$
$A_{3}=$$\left\{y;|y|\geq|x|/2,|x-y|\geq|x|/2\right\}.$
@@
$|v(x,t)|\lesssim t^{-2\alpha}\int_{A_{1}}|f(y)|\mathrm{d}y\lesssim\begin{cases}|X|^{n-2\alpha-\min\{a,n\}}&\mathrm{if~{}~{}}a\neq n\\ |X|^{-2\alpha}\log|X|&\mathrm{if~{}~{}}a=n\end{cases}.$
@@
In $A_{1}$, we have $\frac{|x|}{2}\leq|x-y|\leq\frac{3|x|}{2}$ and $t^{2}+|x-y|^{2}\sim|x|^{2}$, then
@@
$A_{21}=A_{2}\cap\{y;|x-y|>t\}\qquad\mathrm{and}\qquad A_{22}=A_{2}\cap\{y;|x-y|\leq t\}.$
@@
Suppose $f\in L_{\mathrm{loc}}^{1}(\mathbb{R}^{n})$ satisfies that $f(x)=O(|x|^{-b})$ as $|x|\to\infty$ for some $b\in\mathbb{R}_{+}$. Let
$w(x,t)=\int_{\mathbb{R}^{n}}\frac{t^{\beta}f(y)}{(t^{2}+|x-y|^{2})^{\frac{n+\beta}{2}}}\mathrm{d}y,\quad\beta\in\mathbb{R}_{+}.$
@@
$\left|w(x,t)\right|\lesssim$$\begin{cases}|X|^{-\min\{b,n\}}\qquad&\mathrm{if~{}~{}}b\neq n\\ |X|^{-n}\log|X|&\mathrm{if~{}~{}}b=n\end{cases}\qquad\quad\mathrm{~{}~{}as~{}~{}}\quad|X|\to\infty.$
@@
For $r\in\mathbb{R}_{+}$, let $B_{r}^{+}=B_{r}(0)\cap\mathbb{R}^{n+1}_{+}$ and $\partial^{+}B_{r}^{+}=\partial B_{r}^{+}\cap\mathbb{R}^{n+1}_{+}$. Let $P_{k}(x)$ denote a polynomial with degree $\leq k$, where $k\in\mathbb{N}$. Given $u\in C^{\infty}(\overline{\mathbb{R}^{n+1}_{+}})$, we introduce a Kelvin transform for $\Delta^{2}$ by
$u^{\ast}(X)=|X|^{3-n}u\left(X|X|^{-2}\right),\qquad X\in\overline{\mathbb{R}_{+}^{n+1}}\backslash\{0\}.$
For our purpose, we need to a refinement of Sun-Xiong’s results [36, Proposition 3.1 and Theorem 3.2] for homogeneous biharmonic BVPs on $\mathbb{R}_{+}^{n+1}$.
@@
$\begin{cases}\displaystyle~{}\Delta^{2}u=0&\mathrm{~{}~{}in~{}~{}}\quad~{}\mathbb{R}_{+}^{n+1},\\ \displaystyle~{}~{}~{}\partial_{t}u=0&\mathrm{~{}~{}on~{}~{}}\quad\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\partial_{t}\Delta u=0\qquad\qquad&\mathrm{~{}~{}on~{}~{}}\quad\partial\mathbb{R}^{n+1}_{+}.\end{cases}$ (2.1)
@@
For our purpose, the nonnegativity assumption of solutions in Sun-Xiong [36] is redundant.
@@
First, rewrite the boundary conditions as $\mathscr{B}_{1}^{3}u=0,\mathscr{B}_{3}^{3}u=0$. Then the conformal covariance of $\mathscr{B}_{1}^{3}$ and $\mathscr{B}_{3}^{3}$ yields that $u^{\ast}$ satisfies problem (2.1) in $\overline{\mathbb{R}_{+}^{n+1}}\backslash\{0\}$.
@@
Notice that $\eta_{\varepsilon}\in C^{3,1}(\mathbb{R}^{n+1}_{+})$ and $u^{*}$ is biharmonic on +$\overline{\mathbb{R}^{n+1}_{+}}\backslash\{0\}$. Then integrating by parts gives
@@
For $r\in\mathbb{R}_{+}$, we set $v_{r}(X)=u^{*}(rX)$, which is biharmonic in +$\overline{\mathbb{R}^{n+1}_{+}}\backslash\{0\}$. Now we extend $v_{r}$ by an even reflection to $\mathbb{R}^{n+1}\backslash\{0\}$ and apply the local estimates for biharmonic functions (see e.g., [33, Proposition 4]) to obtain
@@
so $\tilde{u}\in C^{4}(\mathbb{R}^{n+1})$ and is biharmonic in $\mathbb{R}^{n+1}$. By [33, Proposition 4] we obtain
@@
If we relax the above (2.2) to the assumption that $u(X)\geq-C|X|^{\delta}$ for some constant $40<\delta<$ 4, which ensures that $u^{\ast}_{-}\in L^{1}(B_{1}^{+})$, then the above argument yields
@@
$\begin{cases}\displaystyle\Delta^{2}u=0&\mathrm{~{}~{}in~{}~{}}\qquad~{}\mathbb{R}_{+}^{n+1},\\ \displaystyle\quad~{}u=0&\mathrm{~{}~{}on~{}~{}}\qquad\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle~{}~{}\Delta u=0\qquad\qquad&\mathrm{~{}~{}on~{}~{}}\qquad\partial\mathbb{R}^{n+1}_{+}.\end{cases}$ (2.4)
@@
$u(X)\geq-C|X|^{-\delta}\quad\mathrm{as}\quad|X|\to\infty.$ (2.5)
@@
Notice that these boundary conditions are equivalent to $\mathscr{B}_{0}^{3}u=0,\mathscr{B}_{3}^{3}u=0$ on $\partial\mathbb{R}_{+}^{n+1}$. By conformal covariance, $u^{\ast}$ satisfies problem (2.4) on +$\overline{\mathbb{R}^{n+1}_{+}}\backslash\{0\}$. Following the same lines as before, we have $u^{\ast}\in L^{1}(B_{1}^{+})$ and $u(x,t)=o(|X|^{4})$ as $|X|\to\infty$. Finally, we extend $u$ to $\mathbb{R}^{n+1}$ by an odd reflection
$\tilde{u}(x,t)=\begin{cases}\displaystyle u(x,t)\qquad&\mathrm{~{}~{}for~{}~{}}\quad t\geq 0,\\ \displaystyle-u(x,-t)\qquad&\mathrm{~{}~{}for~{}~{}}\quad t<0.\end{cases}$
A similar argument as in Lemma 2.3 combined with the boundary conditions shows
@@
If we relax the above assumption (2.5) to $u(X)\geq-C|X|^{\delta}$ as $|X|\to\infty$ for some $30<\delta<$ 3, then the same result in Lemma 2.4 is true.
@@ 3 Biharmonic Poisson kernel
For a smooth compact Riemannian manifold $(X^{n+1},g)$ with boundary $M^{n}=\partial X$ and $n\geq 4$, under the assumption that a certain conformal invariant $\lambda_{1}(P_{4}^{g})>0$ (see [9]), the existence and uniqueness of smooth solutions to the following BVP was studied in [9, Proposition 1.3]: For $\bar{f}_{0},\bar{f}_{1}\in C^{\infty}(M)$,
$\begin{cases}\displaystyle~{}P_{4}^{g}U=0&\mathrm{~{}~{}in~{}~{}}\qquad~{}X,\\ \displaystyle\mathscr{B}^{3}_{0}U=\bar{f}_{0}&\mathrm{~{}~{}on~{}~{}}\qquad M,\\ \displaystyle\mathscr{B}^{3}_{1}U=\bar{f}_{1}\qquad\qquad&\mathrm{~{}~{}on~{}~{}}\qquad M.\end{cases}$
Our purpose is to derive explicit representation formulae of biharmonic Poisson kernels for $\mathbb{R}_{+}^{n+1}$ and $\mathbb{B}^{n+1}$. On $(\mathbb{B}^{n+1},\mathbb{S}^{n},|\mathrm{d}\xi|^{2})$, for each proper conformal boundary operator pair, such a biharmonic BVP is expected to be better understood, serving as a first step to the prescribing boundary curvature problem on manifolds with boundary.
3.1 The half-space
We start with the classical Poisson kernel for $-\Delta$ in $\mathbb{R}_{+}^{n+1}$:
@@ 3.1 The half-space
Suppose $f\in C^{2}(\mathbb{R}^{n})\cap L^{\infty}(\mathbb{R}^{n})$ and $\alpha>1$. Then for every fixed $x\in\mathbb{R}^{n}$, we have
@@
It is not hard to see that the above result is still true provided both $f$ and $|\nabla f|$ are bounded. This elementary lemma will be used later.
@@
Therefore, with the choice of $\sigma$ we combine these estimates to conclude that for $t\ll 1$,
@@
solves $\Delta u=0$ in $\mathbb{R}_{+}^{n+1}$ and $u=f(x)$ on $\partial\mathbb{R}_{+}^{n+1}$. However, for each $x\notin\mathrm{supp}(f)$, the Hopf boundary point lemma yields $\partial_{t}u>0$ on $\partial\mathbb{R}_{+}^{n+1}$.
Proof of Theorem 1.1. It suffices to show that fix $i,k\in\{0,1,2,3\}$, let
@@
Case 1. $(\mathscr{B}^{3}_{0},\mathscr{B}^{3}_{1})$-boundary condition.
@@
Case 2. $(\mathscr{B}^{3}_{0},\mathscr{B}^{3}_{2})$-boundary condition.
@@
$\partial_{t}\Delta|X|^{3-n}=$ $(3-n)\Delta(t|X|^{1-n})=2(n-1)(n-3)t|X|^{-1-n},$
@@
Case 4. $(\mathscr{B}^{3}_{2},\mathscr{B}^{3}_{3})$-boundary condition.
@@
For $n=3$, we only need to point out several significant differences. Observe that
to show
@@
3.2 The unit ball
Our purpose is to give explicit representation formulae of biharmonic Poisson kernels for $\mathbb{B}^{n+1}$ with the help of conformal equivalence between $\mathbb{B}^{n+1}$ and $\mathbb{R}_{+}^{n+1}$.